OpenStudy (anonymous):

Simplify the expression sec(u) csc(u)-sec(u)sin(u) as a single trigonometric function.

OpenStudy (anonymous):

( sec u)(csc u)-(sec u)(sin u)?

OpenStudy (anonymous):

When in doubt, it's always smart to just convert everything to sin's and cos's first.

OpenStudy (anonymous):

(1/cos u)(1 sin u)-(1/cos u)(sin u)

OpenStudy (anonymous):

1/ sin u

OpenStudy (anonymous):

-.-

OpenStudy (anonymous):

(1/ cos u)(1/ sin u)-(1/cos u)(sin u)

OpenStudy (anonymous):

cot=cos/sin?

OpenStudy (anonymous):

do you mean 1/cot u?

OpenStudy (anonymous):

you're on the right track there, luiz, now get everything over a common denominator

OpenStudy (anonymous):

what on the right side?

OpenStudy (anonymous):

sin u/cos u=tan u

OpenStudy (anonymous):

on the right side

OpenStudy (anonymous):

so wait

OpenStudy (anonymous):

you still there bang?

OpenStudy (anonymous):

ok, (1/cos u)(1/ sin u)- (1/cos u)(sin u)

OpenStudy (anonymous):

I got that part

OpenStudy (anonymous):

any tips or hints?

OpenStudy (anonymous):

I got (1/cos u)(1/ sin u)- (1/cos u)(sin u)

OpenStudy (anonymous):

so...

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

you still there?

OpenStudy (anonymous):

hey Happyfeet?

OpenStudy (loser66):

hey, still need help?

OpenStudy (anonymous):

I'm sorry, but I'm still having trouble and yes I still need help lol

OpenStudy (loser66):

I saw many people helped you so I let them do. but at the end.. ha!!!

OpenStudy (anonymous):

I got (1/cos u)(1/ sin u)-(1/cos u)(sin u)

OpenStudy (anonymous):

wat next?

OpenStudy (loser66):

ok, back to the original one. first off factor sec out.

OpenStudy (loser66):

show me your work, please, follow my guide

OpenStudy (anonymous):

so...

OpenStudy (loser66):

factor sec out, what do you have?

OpenStudy (anonymous):

sec u(1 times 1/sin u)- sec u(1 times sin u)

OpenStudy (loser66):

it's not the way to factor \[\large sec (u) csc (u) - sec (u)sin(u) = sec (u) [css (u) - sin (u)]\] get it?

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

so wait

OpenStudy (anonymous):

I got that part

OpenStudy (loser66):

you see it, right? you can step up from there, right?

OpenStudy (anonymous):

so wait, lol

OpenStudy (loser66):

hahaha.. .ok, I am waiting, my friend, take time, manipulate the part [ ]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so

OpenStudy (anonymous):

sec bracket 1-sin squared u/sin u bracket?

OpenStudy (anonymous):

so sec bracket cos squared u/sin u bracket?

OpenStudy (loser66):

\[sec(u) [ csc (u) - sin (u)]\] and replace \(csc (u) = \dfrac{1}{sin (u)}\) into the bracket, you have \[sec(u) [ \frac{1}{sin(u)}-sin(u)]\] still manipulatate the term in the bracket [ ] make them have the same denominator \[sec(u) [ \frac{1 -sin^2(u)}{sin(u)}]\] get this part?

OpenStudy (anonymous):

yes, lord happyfeet

OpenStudy (loser66):

now replace \(1 - sin^2(u) = cos^2(u) to the numerator you have